OpenStudy (anonymous):

17/16x+x/16=6x+1/8+15/16x Please help solve equation

OpenStudy (pfenn1):

Is this equation correct? \[\frac{17}{16x}+\frac{x}{16}=\frac{6x+1}{8}+\frac{15}{16x}\]

OpenStudy (campbell_st):

multiply every term by \[ 16x \] this will eliminate the denominator and give \[17 + x^2 = 2x (6x +1) + 15\] simplify \[ 17 + x^2 = 12x^2 + 2x + 15\] or \[11x^2 + 2x -2 = 0\] use the general quadratic formula to solve \[x = \frac{ -2 \pm \sqrt{92}}{22}\]